# Research on the Application of Computer Vision Based on Deep Learning in Autonomous Driving Technology

### v1

@@ II-A Overview of Autonomous Driving Technology
Autonomous driving technology is built on a complex system architecture that is dedicated to achieving highly integrated and precise maneuvering control. In this system, the environment mapping and perception module is responsible for extracting key information from sensor data, which is collected in real time by multiple sensors around the vehicle, providing the system with a dynamic understanding of the surrounding environment. Subsequently, the self-state estimation module ensures that the vehicle can accurately grasp its own position and status, which is the cornerstone of ensuring operational safety. All information is gathered in the system supervisor, which is a decision-making core responsible for global path planning and motion planning, as well as converting high-level decision-making instructions into local planning. These high-level planning instructions are then passed to execution systems and controllers, which regulate the actual movement of the vehicle, including starting, steering and braking, to achieve smooth and safe navigation. As shown in Figure 1, the entire architecture demonstrates the high degree of collaboration between information flow and control flow in autonomous driving technology, emphasizing the close interaction between precise perception, intelligent decision-making, and fine movements.
@@ II-B Principles Of Computer Vision Systems
A computer vision system is a specially designed integration whose core function is to capture and interpret visual information. As shown in Figure 2, in this system, image acquisition is completed by a sophisticated video camera that can capture high-quality visual data under different lighting conditions. Through the lens, light and images are passed from the target to be measured to the camera sensor, and the captured image is then sent to the computer for further processing. As the center of the system, the computer executes image processing and analysis algorithms, converting image data into useful visual information, such as object detection, classification or three-dimensional reconstruction [8]. In addition, the system also includes input/output interfaces and control mechanisms to ensure that users can interact with the system for command input and result acquisition. At the same time, the control mechanism is responsible for adjusting equipment such as cameras and light sources to adapt to different operating conditions and testing requirements.
@@ III-A Deep Learning-Driven Image Recognition System
During the training phase, a dataset containing thousands of annotated images, including scenes at different times of day, weather conditions, and urban environments, was selected. Objects (vehicles and pedestrians) in images are annotated with precise bounding boxes, and these annotated data are used to train the network for effective feature learning. For the detection of vehicles and pedestrians, the system implements a two-stage detection framework [19]. It first uses selective search to generate potential target candidate areas, then performs feature extraction through CNN, and finally applies support vector machine (SVM) for classification. During this process, the learning rate is set near $1\times 10^{-3}$ to maintain the stability of the training process [20, 21].

### v2

@@ II-A Overview of Autonomous Driving Technology
Autonomous driving technology is built on a complex system architecture that is dedicated to achieving highly integrated and precise maneuvering control. In this system, the environment mapping and perception module is responsible for extracting key information from sensor data, which is collected in real time by multiple sensors around the vehicle, providing the system with a dynamic understanding of the surrounding environment. Subsequently, the self-state estimation module ensures that the vehicle can accurately grasp its own position and status, which is the cornerstone of ensuring operational safety. All information is gathered in the system supervisor, which is a decision-making core responsible for global path planning and motion planning, as well as converting high-level decision-making instructions into local planning. These high-level planning instructions are then passed to execution systems and controllers, which regulate the actual movement of the vehicle, including starting, steering and braking, to achieve smooth and safe navigation. As shown in Fig. 1, the entire architecture demonstrates the high degree of collaboration between information flow and control flow in autonomous driving technology, emphasizing the close interaction between precise perception, intelligent decision-making, and fine movements.
Figure 1: Principle of autonomous driving technology
@@ II-B Principles Of Computer Vision Systems
A computer vision system is a specially designed integration whose core function is to capture and interpret visual information. As shown in Fig. 2, in this system, image acquisition is completed by a sophisticated video camera that can capture high-quality visual data under different lighting conditions. Through the lens, light and images are passed from the target to be measured to the camera sensor, and the captured image is then sent to the computer for further processing. As the center of the system, the computer executes image processing and analysis algorithms, converting image data into useful visual information, such as object detection, classification or three-dimensional reconstruction [8]. In addition, the system also includes input/output interfaces and control mechanisms to ensure that users can interact with the system for command input and result acquisition. At the same time, the control mechanism is responsible for adjusting equipment such as cameras and light sources to adapt to different operating conditions and testing requirements [9].
@@ III-A Deep Learning-Driven Image Recognition System
During the training phase, a dataset containing thousands of annotated images, including scenes at different times of day, weather conditions, and urban environments, was selected. Objects (vehicles and pedestrians) in images are annotated with precise bounding boxes, and these annotated data are used to train the network for effective feature learning. For the detection of vehicles and pedestrians, the system implements a two-stage detection framework [20]. It first uses selective search to generate potential target candidate areas, then performs feature extraction through CNN, and finally applies support vector machine (SVM) for classification. During this process, the learning rate is set near $1\times 10^{-3}$ to maintain the stability of the training process [21, 22, 23].